OpenStudy (ninafilipyeva):

how do you combine like terms? Please help. Example: 3y+2+2y+5=

OpenStudy (priyar):

add all the y terms first

OpenStudy (ninafilipyeva):

2

OpenStudy (priyar):

no..3y+2y=??

OpenStudy (ninafilipyeva):

5 and 2y

OpenStudy (priyar):

no..i'll give u an example 2y+5y=y(2+5)=y(7)=7y

OpenStudy (priyar):

now try 3y+2y

OpenStudy (ninafilipyeva):

I don't understand you, priyar.

pooja195 (pooja195):

3y+2+2y+5 Dogs with the dogs cats with the cats y=dogs cats=numbers 3y+2y=?

OpenStudy (priyar):

add the coeffecients and put y beside them so add 3+5

pooja195 (pooja195):

we are basically saying 3+2=?

OpenStudy (priyar):

*3+2

OpenStudy (ninafilipyeva):

What is a coeffecients?

OpenStudy (ninafilipyeva):

8

pooja195 (pooja195):

No 3+2=?

OpenStudy (ninafilipyeva):

5

pooja195 (pooja195):

Good so 5y 5+2=?

OpenStudy (ninafilipyeva):

7

OpenStudy (priyar):

coeffecient is a number beside the variable... for ex: 3y --> here 3 is the coeffecient and y is the variable

pooja195 (pooja195):

What would our answer be?

OpenStudy (anonymous):

Let y = an apple 3y+2y= How many apples?

OpenStudy (ninafilipyeva):

thanks priyar

OpenStudy (priyar):

welcome

OpenStudy (priyar):

here... 3y+2y=5y got it?

OpenStudy (ninafilipyeva):

I still don't get it

pooja195 (pooja195):

Read this http://www.mathsisfun.com/algebra/like-terms.html

OpenStudy (ninafilipyeva):

Is it five or seen of fifty-seven

pooja195 (pooja195):

3+2=5 5y

OpenStudy (ninafilipyeva):

But, why are we doing 5+2

OpenStudy (priyar):

no...3*y+2*y=(3+2)*y=5*y=5y

OpenStudy (priyar):

we are first taking the variable terms..adding them and then adding this with the constants

pooja195 (pooja195):

We are combining like terms \[\huge~\rm~~\bf~ \color{red}{ (3y+2y)}+\color{purple}{(2+5) }\]

OpenStudy (ninafilipyeva):

thanks both of ya

OpenStudy (ninafilipyeva):

oh. So three plus 2 with y each is five. Then, 2 plus five since there are 2 y.

OpenStudy (priyar):

always when u see such a q first combine the like terms..Do u know what like terms are?

OpenStudy (ninafilipyeva):

no

OpenStudy (priyar):

like terms means similar terms... here : 3y and 2y are similar right?

OpenStudy (ninafilipyeva):

yep

OpenStudy (priyar):

good

OpenStudy (priyar):

how did u say that?

OpenStudy (ninafilipyeva):

They are similar because of the y.

OpenStudy (ninafilipyeva):

right?

OpenStudy (priyar):

Very good!

OpenStudy (priyar):

Now what are the other 2 terms in the Q?

OpenStudy (ninafilipyeva):

2 and 5

OpenStudy (priyar):

these are like terms are not?

OpenStudy (priyar):

*or

OpenStudy (ninafilipyeva):

yes

OpenStudy (priyar):

good how did u say that?

OpenStudy (ninafilipyeva):

say what?

OpenStudy (priyar):

say that 5 and 2 are like terms..

OpenStudy (ninafilipyeva):

they don't have variables

OpenStudy (ninafilipyeva):

they both don't

OpenStudy (priyar):

Very good

OpenStudy (ninafilipyeva):

so what is the answer

OpenStudy (ninafilipyeva):

5y?

OpenStudy (priyar):

u got 5y by combining the variable terms... what about the other two no's?

OpenStudy (ninafilipyeva):

7

OpenStudy (priyar):

good

OpenStudy (ninafilipyeva):

so it is 7+5y=12y.

OpenStudy (priyar):

no... u can add ONLY like terms..7 does not have a y..

OpenStudy (ninafilipyeva):

??????

OpenStudy (priyar):

So the answer is what u got

OpenStudy (ninafilipyeva):

I don't know this part...

OpenStudy (priyar):

its simply 7+5y

OpenStudy (ninafilipyeva):

just 7+5y thanks you helped a lot

OpenStudy (priyar):

welcome! u did a great job too.. you yourself told me the answer!!

OpenStudy (ninafilipyeva):

I just didn't know if I am right!!

OpenStudy (priyar):

now u know how to such q's...just combine the like terms

OpenStudy (ninafilipyeva):

thanks again

OpenStudy (priyar):

welcome